Effect of Stock Split

Copper Grill Restaurant Corporation wholesales ovens and ranges to restaurants throughout the Southwest. Copper Grill Restaurant Corporation, which had 60,000 shares of common stock outstanding, declared a 3-for-1 stock split.

a. What will be the number of shares outstanding after the split?
___ shares

b. If the common stock had a market price of $132 per share before the stock split, what would be an approximate market price per share after the split?
$ ___ per share
Answer #1

a) Number of shares outstanding after split = 60000*3 = 180000 Shares

b) Market price after split = 132/3 = 44 per share